6. a) Calculate the pH of a solution prepared by dissolving 0.0775 mol acetic acid and 0.0460 mol sodium acetate in 1 L of waPlease show a step by step process of how to get Part B. Thank You!
Answer #1

6. a) pH = pka + log [salt] Facid] 0.0460 = - log (1.8x10-5) + log - = - 0.22 4.52 4.74 - ..pH = 4.52 ] to this coas added of